If 0.500 mol of iodine and 0.500 mol of chlorine are initially placed into a 2.00-L reaction vessel at 25°C, find the concentrations of all entities at equilibrium.


I2(g)     +     Cl2(g)    ⇌    2 ICl(g)          K = 81.9 at 25°C


1
Expert's answer
2021-04-11T23:49:58-0400

"I_{2(g) } + Cl_{2(g) }\u21cc 2 ICl_{(g) }"

I: 0.500mol + 0.500mol = 0mol

C: -xmol + -xmol = +xmol

E: 0.500-x mol + 0.500-x = x


"K_c =\dfrac{x²}{(0.500-x)(0.500-x)}"


"81.9 =\dfrac{x²}{(0.500-x)(0.500-x)}"


"81.9 ( 0.250- x +x\u00b2) = x\u00b2"


"20.475 -81.9x+ 81.9x\u00b2 = x\u00b2"


"80.9x\u00b2- 81.9x +20.475 = 0"


x = 0.45 or 0.56, but the value of K is greater than 1, so product conc. > reactant conc.

"\therefore x=0.56M"